Given the following information answer each of the following questions: - Ca=20 c=MPC=3⁄4 Ia =I=20


1
Expert's answer
2022-01-12T08:34:07-0500

The GDP is:

"Y = C + I = 20 + 0.75Y + 20 = 0.75Y + 40,"

0.25Y = 40,

Y = 160.